OpenStudy (anonymous):

Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.) i got 6/2 pi - 6 help me

OpenStudy (anonymous):

\[Radius = 3\sqrt{2}\] Area of circle = \[18 \pi\] Area of one segment = \[1/4 \times (18\pi - 36)\]

OpenStudy (anonymous):

diameter of the circle = sqrt (6^2 + 6^2) = sqrt72 = 6sqrt2 so radius = 3sqrt2 area of sector formed by radii = 1/2 *( 3sqrt2) ^2 * pi/2 area of triangle = 1/2 * (3sqrt2 ^2) * sin pi/2 area of segment = area sector - area triangle

OpenStudy (anonymous):

Ans = 5.13